OpenStudy (anonymous):

If f(x) = x + 5, find f(-9).

OpenStudy (turingtest):

just put -9 where the x is\[f(x)=x+5\]\[f(-9)=(-9)+5=?\]

OpenStudy (anonymous):

-4?

OpenStudy (turingtest):

yep :D